OpenStudy (anonymous):

Starting from equation V (r) = 1 40 p cos  r2, the potential due to dipole find the magnitude of electric field (Er) of a radial component. For what value of θ Er=0?

OpenStudy (anonymous):

Its V(r)=1/4\[\pi \epsilon\]p cos\[\theta\]/r2

OpenStudy (anonymous):

Can you type this out in latex? Go to http://en.wikibooks.org/wiki/LaTeX/Absolute_Beginners. You'll want to read the mathematics section. Latex is the language that all science is reported in (science journals, etc.). I just don't know what equation your trying to write. You can use the equation editor to help you too.

OpenStudy (nikvist):